Find the roots of the following quadratic equations by factorisation: 
 √2 x 2 + 7x + 5√2 = 0 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

√2 x^2 + 7x + 5√2=0
Taking LHS,
=> √2 x^2 + 5x + 2x + 5√2
=> x (√2x + 5) + √2(√2x + 5)
= (√2x + 5)(x + √2)
The roots of this equation, √2 x^2 + 7x + 5√2=0
are the values of x for which (x - 5)(x + 2) = 0
Therefore, √2x + 5 = 0 or x + √2 = 0
=> x = -5/√2 or x = -√2

Thanks